Julie-Annie Corporation is in the mature stage of its corporate life cycle. The firm has a current price of $65.00and Problem its expected dividend, D,, is $4.55. If you require a 12 percent rate of return, then what is the expected price of the firm's common stock in five years?
Answer #1

Price = 01/(ke-g) D1 = Ke = ke-g = growth rate = $65.00 $4.55 12.00% 7.00% 5.00% Price at end of year 5 = $82.96 (65*(1+5%)^5